find the roots using simple fixed point iteration. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

e^x+x=4
1
Expert's answer
2020-10-22T17:51:58-0400
"e^x+x=4"

We know that there is a solution for the equation "e^x+x-4=0" in "[1,2]."


"x=\ln(4-x)"

"x_{i+1}=\ln(4-x_i)"

"\begin{matrix}
 i & x_i \\
 0 & 1\\
 1 & 1.098612 \\
 2 & 1.065189\\
 3 & 1.076643\\
 4 & 1.072733\\
 5 & 1.074069\\
 6 & 1.073613\\
 7 & 1.073769 \\
 8 & 1.073715\\
 9 & 1.073734\\
 10 & 1.073727\\
 11 & 1.073729 \\
 12 & 1.073729\\
 
\end{matrix}"


"\varepsilon =\big|\dfrac{n_{i+1}-n_i}{n_i}\big|\cdot 100\%"

"\varepsilon =\big|\dfrac{1.098612-1}{1}\big|\cdot 100\%=9.8612\%"

"\varepsilon =\big|\dfrac{1.065189-1.098612}{1.098612}\big|\cdot 100\%=3.0423\%"

"\varepsilon =\big|\dfrac{1.076643-1.065189}{1.065189}\big|\cdot 100\%=1.0753\%"

"\varepsilon =\big|\dfrac{1.072733-1.076643}{1.076643}\big|\cdot 100\%=0.3632\%"

"\varepsilon =\big|\dfrac{1.074069-1.072733}{1.072733}\big|\cdot 100\%=0.1245\%"

"\varepsilon =\big|\dfrac{1.073769-1.073713}{1.073713}\big|\cdot 100\%=0.005029\%"

"\varepsilon =\big|\dfrac{1.074069-1.072733}{1.072733}\big|\cdot 100\%=0.1245\%"

"\varepsilon =\big|\dfrac{1.073613-1.074069}{1.074069}\big|\cdot 100\%=0.0425\%"

"\varepsilon =\big|\dfrac{1.073769-1.073613}{1.073613}\big|\cdot 100\%=0.0145\%"

"\varepsilon =\big|\dfrac{1.073715-1.073769}{1.073769}\big|\cdot 100\%=0.0050\%"

"x=1.073613"


The root is "1.074"